Last year Mason Inc. had a total assets turnover of 1.33 and an equity multiplier of 1.75. Its sales were $375,000 and its net income was $10,549. The CFO believes that the company could have operated more efficiently, lowered its costs, and increased its net income by $5,250 without changing its sales, assets, or capital structure. Had it cut costs and increased its net income in this amount, by how much would the ROE have changed? Select the correct answer. a. 2.48% b. 3.26% c. 3.65% d. 2.09% e. 2.87%
Answer #1

Net Income = $10,549

Asset Turnover = Sales/Total Assets = 1.33

Total Assets = $375,000/1.33 = $281,954.88

Equity Multiplier = Total Assets/Total Equity = 1.75

Total Equity = 281954.88/1.75 = $161,117.07

Return on Equity = Net Income/Total Equity = 10549/161117.07

Return on Equity = 6.54%

If there is operation efficiency,

Net Income = 10549 + 5250 = $15,799

Return on Equity = 15799/161117.07 = 9.80%

Change in ROE = 9.80 - 6.54

Change in ROE = 3.26%

Option B is correct